# Efficient Adaptive Joint Significance Test and Sobel-Type

### 20 Mar 2024 v6

@@ Abstract
Mediation analysis is an important statistical tool in many research fields. Its aim is to investigate the mechanism along the causal pathway between an exposure and an outcome. The joint significance test is widely utilized as a prominent statistical approach for examining mediation effects in practical applications. Nevertheless, the limitation of this mediation testing method stems from its conservative Type I error, which reduces its statistical power and imposes certain constraints on its popularity and utility. The joint significance tests for large-scale mediation hypotheses in genome-wide epigenetic research have been extensively investigated, whereas these methods are not applicable to single or small-scale mediation hypotheses. The proposed solution to address this gap is the adaptive joint significance test for one mediator, a novel data-adaptive test for mediation effect that exhibits significant advancements compared to traditional joint significance test. The proposed method is designed to be user-friendly, eliminating the need for complicated procedures. We have derived explicit expressions for size and power, ensuring the theoretical validity of our approach. Furthermore, we extend the proposed adaptive joint significance tests for small-scale mediation hypotheses with family-wise error rate (FWER) control. Additionally, a novel adaptive Sobel-type approach is proposed for the estimation of confidence intervals for the mediation effects, demonstrating significant advancements over conventional Sobel’s confidence intervals in terms of achieving desirable coverage probabilities. Our mediation testing and confidence intervals procedure is evaluated through comprehensive simulations, and compared with numerous existing approaches. Finally, we illustrate the usefulness of our method by analysing three real-world datasets with continuous, binary and time-to-event outcomes, respectively. The method we propose provides a publicly accessible and user-friendly R package, called $\tt AdapMed$, which can be found at https://github.com/zhxmath/AdapMed.
Keywords: Adaptive Sobel’s test; Confidence intervals; Multiple mediators; Small-scale mediation hypotheses; Survival analysis.
@@ 1 Introduction
In the area of mediation analysis, the joint significance test is an important statistical method for investigating the causal mechanism of mediation effects MacKinnon (2008). However, the main shortcoming of this method is due to the conservative type I error of mediation testing MacKinnon et al. (2002), which largely prevents its popularity for practical users. Recently, He et al. (2023) developed a novel adaptive bootstrap testing framework that can be applied to the joint significance test. Furthermore, the confidence interval for the mediation effect is a crucial aspect in mediation analysis, which is highly valuable in comprehending the mediation mechanism. The literature on mediation analysis lacks a substantial number of studies focusing on confidence interval estimation, where the Sobel-type (or normality-based) method is consistently employed to construct these intervals in practical applications. The performance of Sobel-type confidence interval, however, falls short when the mediation effect is zero. To improve the performances of joint significance test and Sobel-type confidence interval, we propose two novel data-adaptive statistical mediation analysis methods with theoretical verification. Compared to He et al. (2023), the main advantages of our proposed method are as follows: First, the adaptive joint significance test is user-friendly without involving complicated resampling procedures. The adaptive bootstrap method proposed by He et al. (2023) necessitates a resampling-based estimation approach, which can be computationally intensive when dealing with large-scale datasets. Second, the adaptive joint significance test is highly applicable to a wide range of outcome types including continuous, binary, and survival outcomes. However, the application of He et al. (2023)’s method to handle survival outcomes remains unclear. Third, the explicit formulation of the coverage probability for the adaptive Sobel-type confidence interval is provided to ensure the validity of our procedure. However, He et al. (2023) did not consider the topic of confidence interval estimation for mediation effects. Here we point out that the method proposed by He et al. (2023) circumvents the need for knowledge of the asymptotic distribution of estimates, making it particularly valuable in situations where deriving the asymptotic properties is challenging.
@@
The remainder of this paper is organized as follows: In Section 2, we review some details about the traditional joint significance test for mediation effects. Then we propose a novel data-adaptive joint significance test for one mediator, together with the explicit expressions of size and power. Meanwhile, an adaptive Sobel test is also introduced, which shows a significant improvement compared to traditional Sobel’s method. In Section 3, we implement the adaptive joint significance test towards small-scale multiple testing with FWER control. Section 4 introduces a data-driven approach to construct a confidence interval for the mediation effect. Section 5 presents some simulation studies to assess the performance of our method. In Section 6, we perform mediation analysis for three real-world datasets with our proposed method. Some concluding remarks are provided in Section 7. All proof details of theorems are presented in the Supplementary Material. Finally, our method offers a publicly available and user-friendly R package, called $\tt AdapMed$, which can be accessed at https://github.com/zhxmath/AdapMed.
@@
2 Adaptive Joint Significance Test
@@
where $P_{\alpha}$ and $P_{\beta}$ are given in (2.3) and (2.4), respectively. Due to the convenience of JS test in practical application, it has been widely adopted by many research fields, such as the social and biomedical sciences. However, the JS test suffers from overly conservative type I error, especially when both $\alpha=0$ and $\beta=0$ MacKinnon et al. (2002). In fact, this is a long-existed and unresolved problem for the Sobel test. To bridge this gap, we propose a novel data-adaptive method for improving the statistical efficiency of the Sobel test, especially focusing on the conservative issue in the case of $\alpha=\beta=0$.
@@
Similar to He et al. (2023), we try to reasonably distinguish $H_{00}$ from $H_{01}$ and $H_{10}$ using a certain criterion. Under $H_{00}$, we have $\lim_{n\rightarrow\infty}\mathbb{P}(|T_{\alpha}|<\lambda_{n},|T_{\beta}|<\lambda_{n}|H_{00})=1$, where $\lambda_{n}=o(\sqrt{n})$ and $\lambda_{n}\rightarrow\infty$ as the sample size $n\rightarrow\infty$. However, $\lim_{n\rightarrow\infty}\mathbb{P}(|T_{\alpha}|<\lambda_{n},|T_{\beta}|<\lambda_{n}|H_{10})=0$ and $\lim_{n\rightarrow\infty}\mathbb{P}(|T_{\alpha}|<\lambda_{n},|T_{\beta}|<\lambda_{n}|H_{01})=0$. In view of these points, we are likely to regard the null hypothesis $H_{0}$ as $H_{00}$ instead of $H_{01}$ or $H_{10}$ when $\max\{|T_{\alpha}|,|T_{\beta}|\}<\lambda_{n}$. The authors of He et al. (2023) employed a similar approach in constructing bootstrap-based testing statistics for evaluating mediation effects, whereas our focus lies in the development of novel testing statistics without the need for re-sampling procedures. We provide a numerical example to illustrate the above-mentioned spirit. Specifically, we generate a series of random samples from the linear mediation models: $Y=0.5+0.5X+\beta M+\epsilon$ and $M=0.5+\alpha X+e$, where $X$, $\epsilon$ and $e$ follow from $N(0,1)$. The resulting $T_{\alpha}$ and $T_{\beta}$ are obtained by the ordinary least square method. In Figure 2, we present the scatter plots of $T_{max}=\max\{|T_{\alpha}|,|T_{\beta}|\}$ and $\lambda_{n}=\sqrt{n}/\log(n)$ with 1000 repetitions, where the sample size is $n=1000$, the parameters $(\alpha,\beta)$ are chosen as $(\alpha,\beta)=(0,0)$, and (0.25, 0), respectively. The observation from Figure 2 suggests that, under the assumption of $\alpha=\beta=0$, there is a high probability that $T_{max}$ is significantly smaller than $\lambda_{n}$ as $n\rightarrow\infty$. However, $T_{max}$ is asymptotically much larger than $\lambda_{n}$ when $\alpha\neq 0$, and the same conclusion also applies to $\beta\neq 0$.
Motivated by the above findings, we propose a novel adaptive joint significance (AJS) test procedure for (2.1), where the p-value is defined as
@@
and $\Phi_{N(0,1)}(\cdot)$ is the cumulative distribution function of $N(0,1)$. Due to the convenience of Sobel test in practical application, it has been widely adopted by many research fields, such as the social and biomedical sciences. However, the Sobel test suffers from overly conservative type I error, especially when both $\alpha=0$ and $\beta=0$ MacKinnon et al. (2002). From Liu et al. (2022), the Sobel statistic $T_{sobel}$ has an asymptotic normal distribution $N(0,1)$ under $H_{01}$ and $H_{10}$, while its asymptotic distribution is $N(0,1/4)$ in the case of $H_{00}$. Therefore, the performance of traditional Sobel test is conservative when $\alpha=0$ and $\beta=0$. In fact, this is a long-existed and unresolved problem for the Sobel test. Similar to the idea of AJS method, we propose a novel adaptive Sobel (ASobel) test procedure for (2.1), and the p-value is defined as
@@ 3 Multiple Mediation Testing with FWER Control
4 Adaptive Sobel-Type Confidence Intervals
@@
The asymptotic distribution of $\hat{\sigma}_{\alpha\beta}^{-1}({\hat{\alpha}\hat{\beta}-\alpha\beta})$ is $N(0,1/4)$ instead of $N(0,1)$ when $\alpha=\beta=0$, as stated in Liu et al. (2022). To amend the issue of ${\rm CI_{Sobel}}$ in the case of $\alpha=\beta=0$, we propose a novel adaptive Sobel-type confidence interval for $\alpha\beta$ as follows,
@@ 7 Concluding Remarks
In this paper, we have proposed an data-adaptive joint significance mediation effects test procedure. The explicit expressions of size and power were derived. We also have extended the AJS for performing small-scale multiple testing with FWER control. An adaptive Sobel-type confidence interval was presented. Some simulations and three real-world examples were used to illustrate the usefulness of our method. The proposed method in practical applications is supported by an R package called ”AdapMed”. The focus of this study was limited to a single mediator or a small number of mediators, and the inclusion of high-dimensional mediators was beyond the scope of this research. The following papers are recommended for those interested in high-dimensional mediation analysis, including Zhang et al. (2016; 2021b; 2024), Luo et al. (2020), Perera et al. (2022) and Tian et al. (2022), among others. In addition, we refer to He et al. (2023)’s adaptive bootstrap method when it becomes challenging to derive the asymptotic properties of estimates for pathway effects.
@@
Efficient Adaptive Joint Significance Test and Sobel-Type Confidence Interval for Mediation Effect
@@ Appendix A Proofs
$\mathbb{P}(\alpha\beta\in{\rm CI_{ASobel}}|\alpha\neq 0,\beta\neq 0)$ (S.20)
@@
With probability approaching one, it follows from (S.20), (S.21) and (S.22) that $\lim_{n\rightarrow\infty}\mathbb{P}(\alpha\beta\in{\rm CI_{ASobel}}|\alpha\neq 0,\beta\neq 0)=1-\delta.$ This ends the proof.

### 07 Apr 2024 v7

# Efficient Adjusted Joint Significance Test and Sobel-Type

@@ Abstract
Mediation analysis is an important statistical tool in many research fields. Its aim is to investigate the mechanism along the causal pathway between an exposure and an outcome. The joint significance test is widely utilized as a prominent statistical approach for examining mediation effects in practical applications. Nevertheless, the limitation of this mediation testing method stems from its conservative Type I error, which reduces its statistical power and imposes certain constraints on its popularity and utility. The joint significance tests for large-scale mediation hypotheses in genome-wide epigenetic research have been extensively investigated, whereas these methods are not applicable to single or small-scale mediation hypotheses. The proposed solution to address this gap is the adjusted joint significance test for one mediator, a novel data-adjusted test for mediation effect that exhibits significant advancements compared to traditional joint significance test. The proposed method is designed to be user-friendly, eliminating the need for complicated procedures. We have derived explicit expressions for size and power, ensuring the theoretical validity of our approach. Furthermore, we extend the adjusted joint significance tests for small-scale mediation hypotheses with family-wise error rate (FWER) control. Additionally, a novel adjusted Sobel-type approach is proposed for the estimation of confidence intervals for the mediation effects, demonstrating significant advancements over conventional Sobel’s confidence intervals in terms of achieving desirable coverage probabilities. Our mediation testing and confidence intervals procedure is evaluated through comprehensive simulations, and compared with numerous existing approaches. Finally, we illustrate the usefulness of our method by analysing three real-world datasets with continuous, binary and time-to-event outcomes, respectively. The method we propose provides a publicly accessible and user-friendly R package, called $\tt AdapMed$, which can be found at https://github.com/zhxmath/AdapMed.
Keywords: Adjusted Sobel’s test; Confidence intervals; Multiple mediators; Small-scale mediation hypotheses; Survival analysis.
@@ 1 Introduction
In the area of mediation analysis, the joint significance test is an important statistical method for investigating the causal mechanism of mediation effects MacKinnon (2008). However, the main shortcoming of this method is due to the conservative type I error of mediation testing MacKinnon et al. (2002), which largely prevents its popularity for practical users. Recently, He et al. (2023) developed a novel adaptive bootstrap testing framework that can be applied to the joint significance test. Furthermore, the confidence interval for the mediation effect is a crucial aspect in mediation analysis, which is highly valuable in comprehending the mediation mechanism. The literature on mediation analysis lacks a substantial number of studies focusing on confidence interval estimation, where the Sobel-type (or normality-based) method is consistently employed to construct these intervals in practical applications. The performance of Sobel-type confidence interval, however, falls short when the mediation effect is zero. To improve the performances of joint significance test and Sobel-type confidence interval, we propose two novel data-adjusted statistical mediation analysis methods with theoretical verification. Specifically, the main advantages of our proposed method are as follows: First, the adjusted joint significance test and adjusted sobel-type confidence interval are two flexible and data-driven methods. Particularly, it is very convenient to implement the two proposed methods from the view of practical application. i.e., our method is user-friendly without involving complicated procedures. Second, compared to the conventional joint significance test, our test method has significantly improved in terms of size and power. The enhanced powers are especially obvious for those relatively weak mediation effects, which are difficult to be recognized as significant mediators by traditional methods. Third, the explicit formulation of the coverage probability for the adjusted Sobel-type confidence interval is provided to ensure the validity of our procedure.
@@
The remainder of this paper is organized as follows: In Section 2, we review some details about the traditional joint significance test for mediation effects. Then we propose a novel data-adjusted joint significance test for one mediator, together with the explicit expressions of size and power. Meanwhile, an adjusted Sobel test is also introduced, which shows a significant improvement compared to traditional Sobel’s method. In Section 3, we implement the adjusted joint significance test towards small-scale multiple testing with FWER control. Section 4 introduces a data-driven approach to construct a confidence interval for the mediation effect. Section 5 presents some simulation studies to assess the performance of our method. In Section 6, we perform mediation analysis for three real-world datasets with our proposed method. Some concluding remarks are provided in Section 7. All proof details of theorems are presented in the Supplementary Material. Finally, our method offers a publicly available and user-friendly R package, called $\tt AdapMed$, which can be accessed at https://github.com/zhxmath/AdapMed.
@@
2 Adjusted Joint Significance Test
@@
where $P_{\alpha}$ and $P_{\beta}$ are given in (2.3) and (2.4), respectively. Due to the convenience of JS test in practical application, it has been widely adopted by many research fields, such as the social and biomedical sciences. However, the JS test suffers from overly conservative type I error, especially when both $\alpha=0$ and $\beta=0$ MacKinnon et al. (2002). In fact, this is a long-existed and unresolved problem for the Sobel test. To bridge this gap, we propose a novel data-adjusted method for improving the statistical efficiency of the Sobel test, especially focusing on the conservative issue in the case of $\alpha=\beta=0$.
@@
Following He et al. (2023), we try to reasonably distinguish $H_{00}$ from $H_{01}$ and $H_{10}$ using a certain criterion. Under $H_{00}$, we have $\lim_{n\rightarrow\infty}\mathbb{P}(|T_{\alpha}|<\lambda_{n},|T_{\beta}|<\lambda_{n}|H_{00})=1$, where $\lambda_{n}=o(\sqrt{n})$ and $\lambda_{n}\rightarrow\infty$ as the sample size $n\rightarrow\infty$. However, $\lim_{n\rightarrow\infty}\mathbb{P}(|T_{\alpha}|<\lambda_{n},|T_{\beta}|<\lambda_{n}|H_{10})=0$ and $\lim_{n\rightarrow\infty}\mathbb{P}(|T_{\alpha}|<\lambda_{n},|T_{\beta}|<\lambda_{n}|H_{01})=0$. In view of these points, we are likely to regard the null hypothesis $H_{0}$ as $H_{00}$ instead of $H_{01}$ or $H_{10}$ when $\max\{|T_{\alpha}|,|T_{\beta}|\}<\lambda_{n}$. Note that He et al. (2023) employed a similar approach in constructing bootstrap-based testing statistics for evaluating mediation effects, whereas our focus lies in the development of novel testing statistics without the need for re-sampling procedures. We provide a numerical example to illustrate the above-mentioned spirit. Specifically, we generate a series of random samples from the linear mediation models: $Y=0.5+0.5X+\beta M+\epsilon$ and $M=0.5+\alpha X+e$, where $X$, $\epsilon$ and $e$ follow from $N(0,1)$. The resulting $T_{\alpha}$ and $T_{\beta}$ are obtained by the ordinary least square method. In Figure 2, we present the scatter plots of $T_{max}=\max\{|T_{\alpha}|,|T_{\beta}|\}$ and $\lambda_{n}=\sqrt{n}/\log(n)$ with 1000 repetitions, where the sample size is $n=1000$, the parameters $(\alpha,\beta)$ are chosen as $(\alpha,\beta)=(0,0)$, and (0.25, 0), respectively. The observation from Figure 2 suggests that, under the assumption of $\alpha=\beta=0$, there is a high probability that $T_{max}$ is significantly smaller than $\lambda_{n}$ as $n\rightarrow\infty$. However, $T_{max}$ is asymptotically much larger than $\lambda_{n}$ when $\alpha\neq 0$, and the same conclusion also applies to $\beta\neq 0$.
Motivated by the above findings, we propose a novel adjusted joint significance (AJS) test procedure for (2.1), where the p-value is defined as
@@
and $\Phi_{N(0,1)}(\cdot)$ is the cumulative distribution function of $N(0,1)$. Due to the convenience of Sobel test in practical application, it has been widely adopted by many research fields, such as the social and biomedical sciences. However, the Sobel test suffers from overly conservative type I error, especially when both $\alpha=0$ and $\beta=0$ MacKinnon et al. (2002). From Liu et al. (2022), the Sobel statistic $T_{sobel}$ has an asymptotic normal distribution $N(0,1)$ under $H_{01}$ and $H_{10}$, while its asymptotic distribution is $N(0,1/4)$ in the case of $H_{00}$. Therefore, the performance of traditional Sobel test is conservative when $\alpha=0$ and $\beta=0$. In fact, this is a long-existed and unresolved problem for the Sobel test. Similar to the idea of AJS method, we propose a novel adjusted Sobel (ASobel) test procedure for (2.1), and the p-value is defined as
@@ 3 Multiple Mediation Testing with FWER Control
4 Adjusted Sobel-Type Confidence Intervals
@@
The asymptotic distribution of $\hat{\sigma}_{\alpha\beta}^{-1}({\hat{\alpha}\hat{\beta}-\alpha\beta})$ is $N(0,1/4)$ instead of $N(0,1)$ when $\alpha=\beta=0$, as stated in Liu et al. (2022). To amend the issue of ${\rm CI_{Sobel}}$ in the case of $\alpha=\beta=0$, we propose a novel adjusted Sobel-type confidence interval for $\alpha\beta$ as follows,
@@ 7 Concluding Remarks
In this paper, we have proposed an data-adjusted joint significance mediation effects test procedure. The explicit expressions of size and power were derived. We also have extended the AJS for performing small-scale multiple testing with FWER control. An adjusted Sobel-type confidence interval was presented. Some simulations and three real-world examples were used to illustrate the usefulness of our method. The proposed method in practical applications is supported by an R package called ”AdapMed”. The focus of this study was limited to a single mediator or a small number of mediators, and the inclusion of high-dimensional mediators was beyond the scope of this research. The following papers are recommended for those interested in high-dimensional mediation analysis, including Zhang et al. (2016; 2021b; 2024), Luo et al. (2020), Perera et al. (2022) and Tian et al. (2022), among others.
@@
Efficient Adjusted Joint Significance Test and Sobel-Type Confidence Interval for Mediation Effect
@@ Appendix A Proofs
$\mathbb{P}(\alpha\beta\in{\rm CI_{ASobel}}|\alpha\neq 0,\beta\neq 0)$
@@
With probability approaching one, it follows from (A), (S.21) and (S.22) that $\lim_{n\rightarrow\infty}\mathbb{P}(\alpha\beta\in{\rm CI_{ASobel}}|\alpha\neq 0,\beta\neq 0)=1-\delta.$ This ends the proof.